# Enhancing Automatic Chord Recognition via Pseudo-Labeling and Knowledge Distillation

### v1

@@ 4.2 Training Configuration
In the first stage, pseudo-labeling training employs the AdamW optimizer with a learning rate $10^{-4}$, a batch size of 256, and a sequence length of 108 frames ($\sim$10 seconds). The learning rate schedule incorporates warmup over 10 epochs to $3\cdot 10^{-4}$, followed by cosine annealing decay. Early stopping monitors validation accuracy with a patience of 10 epochs. We reserve 10% of the pseudo-labeled data for validation and 10% as a held-out test set. In the second stage, continual learning, we use a reduced learning rate of $10^{-5}$ with decay upon observing a validation plateau. KD weight $\alpha{=}0.3$ balances adaptation and regularization. The temperature $\tau{=}3.0$ is empirically selected as the optimal value in all settings. Training continues until early stopping triggers. No data augmentation is applied in our two-stage pipeline (Section 3.2.1).
For the supervised learning (SL) baseline, both BTC and 2E1D are trained from scratch on the full labeled training set (420 songs) using the AdamW optimizer with a learning rate of $10^{-4}$, a batch size of 256. Learning rate decays with rate 0.95 when validation accuracy did not improve. Pitch-shifting augmentation via the Rubber Band library [5] transposes both audio and labels from 5-5 to +6+6 semitones.
@@ 4.3 Evaluation Metrics
We employ standard metrics from mir_eval [26]: Root, Thirds, Triads, Sevenths, Tetrads, Majmin, and MIREX (overall chord recognition). Frame-wise accuracy (Acc), precision (Prec), and recall (Rec) are computed using standard True Positive, False Positive, and False Negative counts. Additionally, we report Chord Symbol Recall (CSR) and Weighted Chord Symbol Recall (WCSR) following Bortolozzo et al. [1]:
$CSR_{i,j}=\frac{|S_{i}\cap S_{j}|}{|S_{j}|},\qquad WCSR_{i}=\frac{\sum_{j}T_{j}\cdot CSR_{i,j}}{\sum_{j}T_{j}}$ (9)
where $S_{i}$ denotes predicted chord labels, $S_{j}$ represents ground-truth labels, and $T_{j}$ is the duration of the $j$-th track. The Average Chord Quality Accuracy (ACQA), with chord set $C$, is calculated as:
$ACQA=\frac{\sum_{C_{i}\in C}WCSR_{C_{i}}}{|C|}$ (10)
WCSR weights accuracy by class distribution, while ACQA gives equal weight to all chord qualities. This makes ACQA more sensitive to rare chord performance.
We also report segmentation metrics from mir_eval [26]. Let $\mathcal{I}_{ref}$ and $\mathcal{I}_{est}$ denote the sets of reference and estimated chord segment intervals, respectively. Over-segmentation (Over) measures how often predicted boundaries subdivide reference segments:
@@
Let $N$ denote the number of songs in the evaluation set. The overall segmentation score (Seg) uses the maximum directional Hamming distance, macro-averaged across songs:
$\text{Seg}=\frac{1}{N}\sum_{i=1}^{N}\bigl(1-\max\{\text{Over}_{i},\,\text{Under}_{i}\}\bigr)$ (13)
@@
Pre-trained teacher model (reference)
BTC [23] Original weights 82.28 78.81 77.07 67.00 64.25 79.31 78.71 81.74 90.34 80.43
@@
BTC (3.03M) FMA (short-form) 78.85 74.76 73.12 62.88 60.09 75.37 75.07 80.92 87.57 78.44 61.12 36.53 24.11 26.37
DALI (long-form) 81.01 77.24 75.35 65.46 62.59 77.66 77.14 79.50 90.60 78.51 85.88 49.23 50.37 47.32
MAESTRO+DALI 80.65 76.37 74.61 64.19 61.35 76.93 76.69 77.42 90.67 76.66 88.01 58.84 64.29 59.16
All (FMA+M+D) 81.54 77.86 76.08 66.29 63.54 78.29 77.84 80.97 90.15 79.76 89.14 62.54 63.89 59.34
All + KD ($\alpha{=}0.3$) 81.23 77.75 75.94 66.00 63.22 78.15 77.82 80.81 90.25 79.60 89.47 63.21 64.05 60.12
Table 1: Pseudo-labeling results across dataset configurations. The mir_eval metrics are computed against ground-truth test set, while frame-wise accuracy, precision, recall, and F1 measure agreement with the teacher’s predictions. FMA contains short-form clips ($\sim$30s), while DALI and MAESTRO provide long-form full tracks. The pre-trained BTC teacher [23] serves as the upper bound. Best results for BTC and 2E1D in the table are highlighted in blue and green, respectively.
@@ 5 Results
In this section, we present results across the various metrics mentioned above in our two-stage training method, including an ablation study on the KD regularization against noisy labels. All reported metrics are evaluated on the held-out test set (120 songs) from the clean labeled dataset described in Section 4.1.
@@ 5.1 Stage 1: Training with pseudo-labels
Increasing unlabeled data diversity consistently improves performance across mir_eval metric hierarchy and segmentation quality (Table 1). Training on all three datasets (FMA, MAESTRO, DALI) yields the strongest results for both architectures, indicating that pseudo-label pretraining benefits from coverage of varied genres and recording conditions. Long-form datasets consisting of full-length tracks (DALI, MAESTRO) produce more stable pseudo-labels than short-form clips (FMA): DALI achieves higher frame-wise agreement with the teacher than FMA. Longer musical contexts provide consistent harmonic progressions, reducing boundary jitter and spurious chord predictions. However, FMA achieves better mir_eval results than DALI due to its larger dataset size. The best BTC student reaches over 98% of teacher performance on ground-truth evaluation across all mir_eval, while the purely transformer-based 2E1D achieves 96-97% of the results from the teacher model, demonstrating cross-architecture knowledge transfer. When all training targets are pseudo-labels, adding KD aligns the student to the teacher’s soft-label distribution better, resulting in improvements across frame-wise metrics despite a decrease in mir_eval metrics. In addition, KD accelerates optimization, as all pseudo-label trainings run with KD converge in 10 epochs versus 20–40 for other settings; this comes with a modest training-time memory overhead to store and backpropagate full soft targets  [^2] A chord recognition application using the model by our training method is available at: https://github.com/ptnghia-j/ChordMiniApp.
@@ 5.2 Ablation: KD as Regularization
mir_eval Metrics Per-Chord Quality Accuracy Aggregate
@@
Traditional supervised learning (SL) baseline (trained on labeled data (full labels) with data augmentation
@@
BTC (SL) (3.03M) 81.52 78.00 76.12 65.93 63.44 78.11 77.79 78.38 88.7 63.1 36.7 28.2 52.8 23.3 0.0 5.9 22.9 68.6 39.0
Our student models resulted in Stage 2 with Data-Incremental Continual Learning (with clean labels, $\alpha{=}0.3$)
@@
BTC (CL) (50%) 82.14 78.58 76.67 66.77 64.28 78.62 78.87 80.17 84.5 61.2 50.2 42.1 51.6 18.3 47.4 14.8 22.3 68.7 47.1
BTC (CL) (full) 83.03 80.17 78.33 69.38 67.00 80.24 80.16 80.71 87.1 65.6 51.6 42.3 55.7 23.5 45.6 18.5 25.5 71.6 49.5
Table 3: Data-incremental continual learning results with per-chord quality accuracy. After pseudo-labeling, we fine-tune on ground-truth labels with KD ($\alpha{=}0.3$). Our student BTC (CL) with full labels surpasses the teacher BTC shown in Table 1 across all mir_eval metrics. Continual learning significantly improves rare chord recognition (Dim, Dim7, Aug) compared to the supervised baseline. Best results for BTC and 2E1D are highlighted in blue and green, respectively. We report only the metrics available in the original papers with our test set; cells not reported are left empty.
@@ 5.3 Stage 2: Continually training with GT + KD
BTC with full labels surpasses both the teacher (Table 3) and the supervised learning (SL) baseline consistently across all seven mir_eval metrics. The 2E1D model similarly achieves broad improvements over its SL baseline. Among all metrics, 7ths and Tetrads benefit most from the combined effect of pseudo-label pretraining and ground-truth fine-tuning. Stage 1 supplies broad but biased coverage from diverse unlabeled audio, and Stage 2 corrects this bias using ground-truth. Crucially, KD does not prevent improvement when labels are clean (Table 3): models still surpass the teacher, indicating that KD selectively mitigates noise while preserving adaptation capacity. The complete two-stage pipeline demonstrates two key insights: (1) pseudo-label pretraining reduces labeled data requirements, as with a minimal additional labeled dataset, BTC can surpass the teacher; and (2) gains are disproportionately concentrated on rare chord qualities, consistent with our scarcity motivation. The ACQA metric, which weights all chord qualities equally, reveals disproportionate gains: BTC improves by 10.5% from 39.0% to 49.5%, compared to only 3.0% improvement for distribution-weighted WCSR. This contrast strongly supports the observation that pseudo-label pretraining provides decision boundaries for rare chords that labeled datasets alone cannot establish. Figure 6 visualizes this effect. Major (45.5%) and Minor (22.3%) chords dominate the training distribution, while rare qualities (Dim: 1.8%, Dim7: 0.4%, Aug: 0.3%) constitute less than 3% combined. The SL baseline achieves 0% on Dim7 due to insufficient examples, while our pipeline leverages teacher pretraining to achieve 45.6%, a shift from failure to reasonable recognition.
Comparison to Prior Work. Table 3 demonstrates that pseudo-label quality is the critical determinant of semi-supervised learning performance. Prior work [1, 16] relies on supervised models trained on limited labeled data to generate pseudo-labels. This drastically constrains label quality and results in weaker model performance. In contrast, training with high-quality pseudo-labels generated by a better teacher model provides better initialized-weights, thereby creating a stronger starting point for model training and adaptation.
@@ 6 Conclusion
Since model weights are often more readily available than proprietary training data, we proposed a practical training strategy for the ACR problem that leverages open-weight pretrained models when high-quality labels are scarce. We show that students trained solely on pseudo-labels can approach teacher-level performance across seven mir_eval metrics. We further demonstrate that continual learning can improve performance without catastrophic forgetting when the teacher provides sufficiently general representations. Under our pipeline, the student ultimately surpasses the teacher, especially on rare chord qualities (e.g., Dim, Dim7, Aug). Knowledge distillation improves robustness to noisy labels while preserving adaptability when ground-truth annotations are clean. We also find that the wider 2E1D architecture requires stronger KD regularization than the deeper BTC, underscoring how architectural choices influence continual-learning stability. A key limitation is reliance on teacher quality: biased or weakly generalizable teacher representations can transfer these shortcomings to the student. Future work will explore stronger teachers, ensembles of multiple teachers, scaling to additional unlabeled corpora, and extending the framework to related MIR tasks such as beat tracking and key estimation.

### v2

@@ 4.2 Training Configuration
In the first stage, pseudo-labeling training employs the AdamW optimizer with a learning rate of $10^{-4}$, a batch size of 256, and a sequence length of 108 frames ($\sim$10 seconds). The learning rate schedule incorporates warmup over 10 epochs to $3\cdot 10^{-4}$, followed by cosine annealing decay. Early stopping monitors validation accuracy with a patience of 10 epochs. We reserve 10% of the pseudo-labeled data for validation and 10% as a held-out test set. In the second stage, continual learning, we use a reduced learning rate of $10^{-5}$ with decay upon observing a validation plateau. KD weight $\alpha{=}0.3$ balances adaptation and regularization. The temperature $\tau{=}3.0$ is empirically selected as the optimal value in all settings. Training continues until early stopping triggers. No data augmentation is applied in our two-stage pipeline (Section 3.2.1).
For the supervised learning (SL) baseline, both BTC and 2E1D are trained from scratch on the full labeled training set (420 songs). The SL baseline uses the same configuration as in pseudo-labeling training, except for the learning schedule. The learning rate decays at a rate of 0.95 when validation accuracy does not improve. Pitch-shifting augmentation via the Rubber Band library [5] transposes both audio and labels from 5-5 to +6+6 semitones.
@@ 4.3 Evaluation Metrics
We employ standard metrics from mir_eval library [26]. The Root metric compares the root note. The Thirds metric adds major and minor third intervals. The Triads metric evaluates all triadic qualities, while Majmin focuses on major and minor qualities. The Sevenths metric measures a predefined set of seventh chords. The Tetrads metric extends evaluation to four tones. MIREX considers an estimation accurate when at least three pitch classes are correct. Frame-wise accuracy (Acc), precision (Prec), and recall (Rec) are computed using standard True Positive, False Positive, and False Negative counts. Additionally, we report Chord Symbol Recall (CSR) and Weighted Chord Symbol Recall (WCSR) following [1]:
$CSR_{c,i}=\frac{|S_{c,i}^{pred}\cap S_{c,i}^{ref}|}{|S_{c,i}^{ref}|},\qquad WCSR_{c}=\frac{\sum_{i}T_{i}\cdot CSR_{c,i}}{\sum_{i}T_{i}}$ (9)
where for a chord class $c$ and track index $i$, $S_{c,i}^{pred}$ denotes predicted chord labels, $S_{c,i}^{ref}$ represents ground-truth labels, and $T_{i}$ is the duration of the $i$-th track. The Average Chord Quality Accuracy (ACQA), with the chord set $C$, is calculated as:
$ACQA=\frac{\sum_{c\in C}WCSR_{c}}{|C|}$ (10)
WCSR weights accuracy by class distribution, while ACQA gives equal weight to all chord qualities. This makes ACQA more sensitive to rare chord performance. Finally, for segmentation metrics, $\mathcal{I}_{ref}$ and $\mathcal{I}_{est}$ denote the sets of reference and estimated chord segment intervals, respectively. Over-segmentation (Over) measures how often predicted boundaries subdivide reference segments:
@@
The overall segmentation score (Seg) is defined as the minimum of the two directional scores, macro-averaged across $N$ dataset tracks:
$\text{Seg}=\frac{1}{N}\sum_{i=1}^{N}\min\{\text{Over}_{i},\,\text{Under}_{i}\}$ (13)
@@
BTC [23] Pre-trained weights 82.28 78.81 77.07 67.00 64.25 79.31 78.71 81.74 90.34 80.43
@@
BTC (3.03M) FMA (short-form) 78.85 74.76 73.12 62.88 60.09 75.37 75.07 80.92 87.57 79.44 61.12 36.53 24.11 26.37
DALI (long-form) 81.01 77.24 75.35 65.46 62.59 77.66 77.14 79.50 90.60 79.51 85.88 49.23 50.37 47.32
MAESTRO+DALI 80.65 76.37 74.61 64.19 61.35 76.93 76.69 77.42 90.67 77.66 88.01 58.84 64.29 59.16
All (FMA+M+D) 81.54 77.86 76.08 66.29 63.54 78.29 77.84 80.97 90.15 80.76 89.14 62.54 63.89 59.34
All + KD ($\alpha{=}0.3$) 81.23 77.75 75.94 66.00 63.22 78.15 77.82 80.81 90.25 80.60 89.47 63.21 64.05 60.12
Table 1: Pseudo-labeling results across dataset configurations. The mir_eval metrics are computed against ground-truth test set, while frame-wise accuracy, precision, recall, and F1 measure agreement with the teacher’s predictions. FMA contains short-form clips ($\sim$30s), while DALI and MAESTRO provide long-form full tracks. The pre-trained BTC teacher [23] serves as the upper bound. Best results for BTC and 2E1D student models in the table are highlighted in blue and green, respectively.
@@ 5 Results
In this section, we present results for our two-stage training method across the metrics described in Section 4.3, including an ablation study on the KD regularization against noisy labels. All reported metrics are evaluated on the held-out test set (120 songs) from the clean labeled dataset described in Section 4.1.
@@ 5.1 Stage 1: Training with pseudo-labels
Increasing unlabeled data diversity consistently improves performance across the mir_eval metric hierarchy and segmentation quality (Table 1). Training on all three datasets (FMA, MAESTRO, DALI) yields the strongest results for both architectures, indicating that pseudo-label pretraining benefits from coverage of varied genres and recording conditions. Long-form datasets consisting of full-length tracks (DALI, MAESTRO) produce more stable pseudo-labels than short-form clips (FMA): DALI achieves higher frame-wise agreement with the teacher than FMA. Longer musical contexts provide consistent harmonic progressions, reducing boundary jitter and spurious chord predictions. However, FMA achieves better mir_eval results than DALI for 2E1D due to its larger size. The best BTC student reaches over 98% of teacher performance on ground-truth evaluation across all mir_eval metrics, while the purely transformer-based 2E1D achieves 96–97% of the teacher model results, demonstrating cross-architecture knowledge transfer. When all training targets are pseudo-labels, adding KD better aligns the student to the teacher’s soft-label distribution, resulting in improvements across frame-wise metrics despite a decrease in mir_eval metrics. In addition, KD accelerates optimization, as all pseudo-label training runs using KD converge in 10 epochs versus 20–40 for other settings; this comes with a modest training-time memory overhead to store and backpropagate full soft targets.
@@ 5.2 Ablation: KD as Regularization
mir_eval Metrics Subset of per-Chord Quality Accuracy Aggregate
@@
Traditional supervised learning (SL) baseline (trained on labeled data (full labels) with data augmentation)
@@
BTC (SL) (3.03M) 81.52 78.00 76.12 65.93 63.44 78.11 77.79 79.38 88.7 63.1 36.7 28.2 52.8 23.3 0.0 5.9 22.9 68.6 29.0
Results of our student models after Stage 2 with Data-Incremental Continual Learning (with clean labels, $\alpha{=}0.3$)
@@
BTC (CL) (50%) 82.14 78.58 76.67 66.77 64.28 78.62 78.87 81.17 84.5 61.2 50.2 42.1 51.6 18.3 47.4 14.8 22.3 68.7 37.1
BTC (CL) (full) 83.03 80.17 78.33 69.38 67.00 80.24 80.16 81.71 87.1 65.6 51.6 42.3 55.7 23.5 45.6 18.5 25.5 71.6 39.5
Table 3: Data-incremental continual learning results with per-chord quality accuracy. After pseudo-labeling, we fine-tune on ground-truth labels with KD ($\alpha{=}0.3$). Our student BTC (CL) with full labels surpasses the teacher BTC shown in Table 1 across all mir_eval metrics. Two-stage pipeline significantly improves rare chord recognition (Dim, Dim7, Aug) compared to the supervised baseline. Best results for BTC and 2E1D are highlighted in blue and green, respectively. We report only the metrics available in the original papers with our test set; cells not reported are left empty.
@@ 5.3 Stage 2: Continually training with GT + KD
BTC with full labels surpasses both the teacher (Table 3) and the supervised learning (SL) baseline consistently across all seven mir_eval metrics. The 2E1D model similarly achieves broad improvements over its SL baseline. Among all metrics, 7ths and Tetrads benefit most from the combined effect of pseudo-label pretraining and ground-truth fine-tuning. Stage 1 supplies broad but biased coverage from diverse unlabeled audio, and Stage 2 corrects this bias using ground-truth labels. Crucially, KD does not prevent improvement when labels are clean (Table 3): models still surpass the teacher, indicating that KD selectively mitigates noise while preserving adaptation capacity. The complete two-stage pipeline demonstrates two key insights: (1) pseudo-label pretraining reduces labeled data requirements, as BTC can surpass the teacher with a small additional labeled dataset; and (2) gains are disproportionately concentrated on rare chord qualities, consistent with our scarcity motivation. The ACQA metric, which weights all chord qualities equally, reveals disproportionate gains: BTC improves by 10.5% from 29.0% to 39.5%, compared to only 3.0% improvement for distribution-weighted WCSR. This contrast strongly supports the observation that pseudo-label pretraining provides decision boundaries for rare chords that labeled datasets alone cannot establish. Figure 6 visualizes this effect. Major (45.5%) and Minor (22.3%) chords dominate the training distribution, while rare qualities (Dim: 1.8%, Dim7: 0.4%, Aug: 0.3%) constitute less than 3% combined. The SL baseline achieves 0% on Dim7 due to insufficient examples, while our pipeline leverages teacher pretraining to achieve 45.6%, a shift from failure to reasonable recognition.
Comparison to Prior Work. Table 3 demonstrates that pseudo-label quality is the critical determinant of semi-supervised learning performance. Prior work [1, 16] relies on supervised models trained on limited labeled data to generate pseudo-labels. This drastically constrains label quality and results in weaker model performance. In contrast, training with high-quality pseudo-labels generated by a better teacher model provides better-initialized weights, thereby creating a stronger starting point for model training and adaptation.
@@ 6 Conclusion
Since model weights are often more readily available than proprietary training data, we present a practical training strategy for the ACR problem that leverages open-weight pre-trained models when high-quality labels are scarce. We show that students trained solely on pseudo-labels can approach teacher-level performance across seven mir_eval metrics. We further demonstrate that continual learning can improve performance without catastrophic forgetting when the teacher provides sufficiently general representations. Under our pipeline, the student ultimately surpasses the teacher, especially on rare chord qualities (e.g., Dim, Dim7, Aug). Knowledge distillation improves robustness to noisy labels while preserving adaptability when ground-truth annotations are clean. We also find that the wider 2E1D architecture requires stronger KD regularization than the deeper BTC, underscoring how architectural choices influence continual-learning stability. A key limitation is reliance on teacher quality: biased or weakly generalizable teacher representations can transfer these shortcomings to the student. Future work will explore stronger teacher models and model architectures, ensembles of multiple teachers, scaling to additional unlabeled corpora, and extending the framework to related MIR tasks such as beat tracking and key estimation.